Katy:

In a circle, a 90° sector has area 36π square inches. What is the radius of the circle? 36 inches 18 inches 12 inches 6 inches

Hero:

Can I help?

Katy:

Yes, please

Katy:

Can you help me solve that question?

Hero:

Okay, so here is a circle:  Now here is a circle with a \(90^{\circ}\) sector:  How much of the circle does that \(90^{\circ}\) cover?

Katy:

A fourth of that circle

Hero:

Correct

Katy:

1/4 also 25% as a percentage

Hero:

Correct, so if we have the area of one-fourth of the circle, what should we do to get the area of the whole circle?

Katy:

36 divided by 4 times 3.14?

Hero:

Put it this way: If I have \(\dfrac{1}{4}\) of something, what do I need to multiply \(\dfrac{1}{4}\) by to get \(1\) ?

Katy:

9 times 3.14?

Katy:

Probably not 9 times 3.14, I don't understand.

Hero:

\(\dfrac{1}{4} \times 4 = \) ?

Katy:

Oh, 1/4*4=1

Hero:

Exactly and they gave you one fourth the area of the circle.

Katy:

Is it 36 divided by 4?

Hero:

What operation did you just perform? Division or multiplication?

Katy:

1/4*4=1 would be multiplication

Hero:

Correct. You had one-fourth and you had to multiply by 4 to get 1. Now you have one fourth of the area of a circle and you need to get the area of one whole circle. So the operation must remain consistent.

Katy:

Do I multiple 1 by something?

Hero:

When you had one-fourth, what did you multiply by to get one?

Katy:

Do I multiply 4 by something?

Hero:

Yes

Katy:

Is it 4 times 3.14?

Hero:

What was the given area?

Hero:

\(36\pi\) right?

Katy:

Do I multiply 4 by 36?

Hero:

Yes, by \(36\pi\)

Hero:

We figured out that the \(36\pi\) was one-fourth of the circle. In order to get the area of the whole circle, we multiply by four.

Katy:

36 times 3.14 is 113.04?

Katy:

4 times 113.04?

Hero:

@Ultrilliam we need a calculator on here

Katy:

Is it 4 times 113.04?

Ultrilliam:

Fine, I'll get to work on one, it shouldn't be too difficult to add

Hero:

Thank you

Hero:

Yes Katy, but we need to write it as \(36 \times 4 \times \pi\)

Hero:

Because we want to find the exact radius

Katy:

Okay

Hero:

What is \(36 \times 4\)?

Katy:

452.16

Katy:

36 times 4 is different

Hero:

Because we don't need to multiply the 3.14 to find the radius

Katy:

144

Hero:

Except we have to write it as \(144\pi\)

Katy:

144 times 3.14 is

Hero:

But we're not done yet

Hero:

What you don't understand is we don't need to multiply 144 by 3.14

Hero:

in order to find the radius

Hero:

What we have to do is find the square root of 144

Katy:

452.16

Hero:

No, we don't need to do that.

Hero:

find the square root of 144

Katy:

Oh

Katy:

12

Hero:

Finally. That is the radius of the circle.

Hero:

It wasn't as difficult as you were making it.

Hero:

You focused too much on multiplying by 3.14 when that step was not necessary to find the radius.

Hero:

They gave us \(36\pi\) which is one fourth the area of the circle.

Hero:

To find the area of the whole circle we multiply by four to get \(144 \pi\) square inches.

Katy:

I have 4 more questions that I will have to do later.

Hero:

Then we must realize we need to re-write the expression in the form \(r^2 \pi\) since this is the same as \(\pi r^2\) which is the area of the circle formula.

Hero:

So since \(144 = 12^2\) we know that the area of the circle can also be written as \(12^2 \pi\) and 12 is the radius in inches.